2. Compute an orthonormal basis for the null space of the linear map T: F – F given by T(a,b,c) = a + b + c.
Answer #1

T: F at btc -> F T9,b,c) = need to find Tra,b,c) = we basis such that O at btc a = b C let b=t c=5 -t-s a = a - t ts o b C S